OpenStudy (agentjamesbond007):

[find limit] limit as x approaches infinity (√(9x^2+x)-3x)

OpenStudy (zarkon):

1/6

OpenStudy (anonymous):

method is to multiply by \[\frac{\sqrt{9x^2+3}+3x}{\sqrt{9x^2+3}+3x}\] unless there is a snap way i don't know zarkon?

OpenStudy (zarkon):

that is the way I did it

OpenStudy (agentjamesbond007):

is it multiplying by the conjugate or how did you get that?